Eastern Polymers, Inc., processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 5,200 units of product were as follows:

Standard Costs Actual Costs
Direct materials 6,800 lbs. at $4.8 6,700 lbs. at $4.7
Direct labor 1,300 hrs. at $16.1 1,330 hrs. at $16.4
Factory overhead Rates per direct labor hr.,
based on 100% of normal
capacity of 1,360 direct
labor hrs.:
Variable cost, $4.8 $6,180 variable cost
Fixed cost, $7.6 $10,336 fixed cost

Each unit requires 0.25 hour of direct labor.

Required:

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Price variance $
Quantity variance $
Total direct materials cost variance $

b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Rate variance $
Time variance $
Total direct labor cost variance $

c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Variable factory overhead controllable variance $
Fixed factory overhead volume variance $
Total factory overhead cost variance $
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
6700 $4.70 $31,490.00 6700 $4.80 $32,160.00 6800 $4.80 $32,640.00
$670.00 F $480.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $670.00 F
Direct material quantity variance $480.00 F
Direct material cost variance $1,150.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
1330 $16.40 $21,812.00 1330 $16.10 $21,413.00 1300 $16.10 $20,930.00
$399.00 U $483.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $399.00 U
Direct Labor Efficiency variance $483.00 U
Direct labor cost variance $882.00 U

Solution c:

Variable factory overhead controllable variance = Standard variable factory overhead cost - Actual variable factory overhead cost

= (1300*$4.8) - $6,180 = $60 F

Fixed overhead volume variance = Fixed overhead applied - budgeted fixed overhead

= (1300*$7.60) - (1360*$7.60) = $456 U

Total factory overhead cost variance = Variable factory overhead controllable variance + Fixed overhead volume variance

= $60 F + $456 U = $396 U